# MeanSparse: Post-Training Robustness Enhancement Through Mean-Centered Feature Sparsification

### v2

@@ 4 Experiments
4.1 Evaluation Metrics
[…]
4.2 Integration to State-of-the-Art Models
Figure LABEL:fig:spaf shows that integrating MeanSparse into the SOTA robust model in the RobustBench ranking [10] across various benchmark datasets achieves a new SOTA in robustness. Figure LABEL:fig:bench:sub1 depicts the results for the Clean and AA metrics of the standard and MeanSparse-integrated versions of the WideResNet-94-16 [3], RaWideResNet-70-16 [34] and WideResNet-70-16 [48] over CIFAR-10 dataset. The original WideResNet-94-16 [3] currently holds the SOTA performance in the RobustBench ranking [10] for $\ell_{\infty}$ untargeted attack. After post-processing with the proposed pipeline, the Clean accuracy drops slightly (from $93.68\%$ to $93.63\%$), while the AA significantly increases (from $73.71\%$ to $75.28\%$), setting a new record for robustness in terms of AA on the CIFAR-10 dataset. Similar results can be obtained for RaWideResNet-70-16 [34] and WideResNet-70-16 [48]. Also, as a result of post-processing the robust model based on the WideResNet-70-16 architecture, the resulting AA accuracy ($71.41\%$) exceeds this value for the robust model based on the RaWideResNet architecture ($71.07\%$), even though WideResNet has a less complex architecture compared to RaWideResNet, which is equipped with the Squeeze-and-Excitation module.
Figure LABEL:fig:bench:sub2 represents the result of integrating MeanSparse into the current SOTA robust model (WideResNet-70-16 [48]) in terms of AA accuracy with $\ell_{2}$ distance metric. Similarly, while the clean accuracy is almost unchanged (from $95.54\%$ to $95.49\%$), the AA accuracy is improved (from $84.97\%$ to $87.28\%$).
Figure LABEL:fig:bench:sub3 illustrates the clean and AA accuracy for three different models adversarially trained on the ImageNet dataset (Swin-L [27], ConvNeXt-L [26] and RaWideResNet [34]). By carefully selecting the $\alpha$ value, while the Clean accuracy remains unchanged, the AA accuracy improves significantly. The Swin-L model [27] is currently the SOTA robust model over the ImageNet dataset. When integrating this model with MeanSparse, while clean accuracy drops slightly (from $78.92\%$ to $78.86\%$), the AA accuracy is increased (from $59.56\%$ to $62.12\%$) setting a new SOTA.
Figure LABEL:fig:bench:sub4 shows the results of integrating MeanSparse with the current state-of-the-art (SOTA) robust model on the CIFAR-100 dataset (WideResNet-70-16 [48]). For this model, the clean and AutoAttack (AA) accuracies are $75.22\%$ and $42.67\%$, respectively. After integrating MeanSparse, these accuracies change to $75.17\%$ and $44.78\%$, establishing a new SOTA robust model with nearly identical clean accuracy.
Table 1 presents the results of integrating MeanSparse into RaWideResNet [34] and WideResNet [48] architectures trained on the CIFAR-10 dataset and ConvNeXt-L [26] and RaWideResNet [34] on the ImageNet dataset using adversarial training. The results are provided for different values of $\alpha$ (note that the threshold equals $\alpha\times\sigma_{\mathrm{ch}}$). The experiment results over the CIFAR-10 test set depict that for a range of $\alpha$ below $0.25$, the clean accuracy has decreased negligibly or almost unchanged while $\text{APGD}_{\text{ce}}$is increasing. Although these results may be attributed to overfitting, the result over the CIFAR-10 training set also represents a similar trend. So an important takeaway from CIFAR-10 results in Table 1 is that the variations around the mean value of representation can be considered as a capacity provided by the model for the attackers while its utilization is limited. MeanSparse requires one pass over CIFAR-10 training set to calculate the statistics. The time for this pass is 102 seconds and 95 seconds for the RaWideResNet and WideResNet architectures, respectively. The increase in $\text{APGD}_{\text{ce}}$evaluation time after MeanSparse integration is negligible, with only a $2\%$ increase in time.
Results over ImageNet presented in Table 1 show that similar outcomes are achieved when MeanSparse is integrated into models based on the ConvNeXt-L [26] and RaWideResNet [34] architectures adversarially trained. Note that throughout our experiment over the RaWideResNet [34] architecture for both CIFAR-10 and ImageNet-1k datasets, we do not apply the sparsification operator inside the Squeeze-and-Excitation module. The time for the MeanSparse pass over the ImageNet-1K training set is 89 minutes for the ConvNeXt-L architecture and 101 minutes for the RaWideResNet architecture. The increase in $\text{APGD}_{\text{ce}}$evaluation time after MeanSparse integration is also negligible, with only a $3\%$ increase in time.
@@ 4.2 Integration to State-of-the-Art Models
Table 1: Clean accuracy and accuracy under A-PGD-CE attack [10] (represented by $\text{APGD}_{\text{ce}}$) for leading robust models on the CIFAR-10 and ImageNet-1k datasets (Base row) and after their integration with the MeanSparse technique (RaWideResNet [34] and WideResNet [48] are leading models in Robustbench ranking for CIFAR-10 dataset and ConvNeXt-L [26] and RaWideResNet [34] are highly ranked models for ImageNet dataset in Robustbench [9]). The best result in each column is in bold and the threshold selected through the proposed pipeline is in blue.
@@
4.3 Insights from Ablation Study
We conducted several experiments to explore different aspects of MeanSparse integration into trained models. These experiments examine the effects of activation functions A.1, the type of adversarial attack (either $\ell_{\infty}$ or $\ell_{2}$) A.2, the robustness of MeanSparse integrated models against black-box attacks A.3, the effect of adaptive attack with full access to MeanSparse modules location and specifications (mean and variance vectors) A.4, the types of adversarial examples used during training with AT A.5, the effect of integrating a standard trained model with MeanSparse A.6, the impact of centering MeanSparse around the mean value A.7, the influence of attack power on MeanSparse integration performance A.8, the effects of integrating MeanSparse with different patterns, such as isolated integration with a single activation A.9, and the robustness improvement resulting from MeanSparse integration in terms of PGD-50 attack A.10. The main takeaways from our ablation study can be summarized as:
  • MeanSparse effectively enhances robustness across various activation functions. In other words, the activation function alone does not effectively eliminate non-robust features, which are previously reduced during training.
  • By calculating the sparsity using other well-known adversarial attacks such as PGD-50 [29], we observe a similar improvement in robustness, consistent with the AA and APGD accuracy values.
  • Although MeanSparse masks the gradient, which can explain the robustness improvement against white-box attacks, it also demonstrates enhanced robustness when evaluated using black-box attacks.
  • Designing adaptive adversarial attacks for models integrated with MeanSparse can slightly reduce the robust accuracy. However, due to the significant improvement from MeanSparse integration, the adaptively attacked models still exhibit substantially higher robustness compared to the original base model (see Section A.4 for a detailed description of the adaptive attack).
  • MeanSparse is effective against both $\ell_{\infty}$ and $\ell_{2}$ attacks. Specifically, if a model is fortified against $\ell_{\infty}$ attacks using AT, adding MeanSparse enhances robustness against both $\ell_{\infty}$ and $\ell_{2}$ attacks. Also, MeanSparse is not sensitive to the method used to generate adversarial samples during AT.
  • Using dedicated mean and variance for each channel in the MeanSparse technique is critical to its performance. Otherwise, we cannot expect the observed improvement in robustness.
  • Integrating MeanSparse in isolated activation functions remains effective for robustification, although the improvement is limited.
4.4 Limitations
The MeanSparse technique proposed in this paper is a post-training method that is useful on top of adversarially trained models. Using MeanSparse for models trained in a standard way does not lead to robustness improvement as we expect based on our intuition introduced in this paper.
5 Conclusions
Adversarial training, a widely accepted defense against evasion attacks, attenuates non-robust feature extractors in a model. Although adversarial training (AT) is effective, this paper demonstrates that there is still an easily accessible capacity for attackers to exploit. We introduce the MeanSparse technique, which partially blocks this capacity in the model available to attackers and enhances adversarial robustness. This technique, easily integrable after training, improves robustness without compromising clean accuracy. By integrating this technique into trained models, we achieve a new record in adversarial robustness on CIFAR-10 for both $\ell_{\infty}$ and $\ell_{2}$ untargeted attacks, as well as on CIFAR-100 and ImageNet for $\ell_{\infty}$ untargeted attacks.
Ethical Considerations
This research is conducted with adherence to the highest ethical standards, ensuring compliance with all relevant institutional and international guidelines to the best of the authors’ knowledge. We critically assessed the societal impacts of our work, aiming to mitigate any negative consequences while promoting beneficial applications. Positive impacts include improved model robustness and increased trust in AI systems. Negative impacts involve potential misuse for malicious applications and reinforcing existing biases. To address these, we suggest mitigation strategies such as implementing gated releases, conducting fairness audits, and promoting transparency through public availability of code and data. Transparency and accountability were prioritized throughout the research process, and there were no conflicts of interest influencing the outcomes of this study.
Appendix A Ablation Experiment
We conducted several experiments to evaluate the effectiveness of MeanSparse in various scenarios. In the majority of experiments, we use APGD accuracy (accuracy after attacking the model with Auto-PGD using both Cross-entropy and Difference of Logits Ratio loss metrics) as the robustness metric [10]. During our simulations, we found APGD accuracy to be almost equal to AA accuracy, with the advantage that APGD can be measured faster.
A.1 Activation Function
MeanSparse is generally appended before the activations function. This questions the effectiveness of MeanSparse for different activation functions. In this experiment, we check the performance of MeanSparse when integrating the models with different activation functions. The specifications of this experiment are:
  • Dataset: CIFAR-10 [23]
  • Architecture: ResNet-18 [20]
  • Optimizer: SGD (initial learning rate: 0.1, weight decay: 0.0005, momentum: 0.9)
  • Number of Epochs: 200
  • Batch Size: 256
  • Learning rate scheduler: The initial learning rate of 0.1 is reduced by a factor of 10 at epoch 100 and 150.
  • Best model selection: We evaluate the model at each epoch of training and select the one with the highest PGD adversarial accuracy on the test set.
  • Adversarial training properties: 10 step PGD adversarial training [29] with respect to $\ell_{\infty}$ attacks with radius $8/255$ and step size of 0.0078.
@@ A.1 Activation Function
Table 2 demonstrates the Clean and APGD accuracy metrics for a base ResNet-18 model, denoted by Base, with different activation functions and the same model post-processed by the proposed pipeline to integrate MeanSaprse. We use both non-parametric and parametric activation functions, such as PSSiLU [11]. In our experiment, the GELU activation function, proposed in [21] and generally used in attention-based architectures, results in the best APGD accuracy. When this model is integrated with MeanSparse, we observe a significant improvement in APGD accuracy ($49.12\%\rightarrow 50.37\%$) while the clean accuracy slightly decreases ($84.59\%\rightarrow 84.32\%$) for $\alpha=0.2$. Additionally, if we can tolerate $\sim 1\%$ reduction in clean accuracy, we can increase APGD accuracy by more than $2\%$. The first important takeaway is the effectiveness of MeanSparse integration with different activations and the inability of the activation function to replicate the role of the MeanSparse technique. The second is that $\alpha=0.2$ consistently leads to large improvements in APDG accuracy with a negligible reduction in clean accuracy for all activation functions. Even for the ReLU activation function, $\alpha=0.2$ increases both Clean and APGD accuracies simultaneously. We will use the ResNet-18 model trained adversarially with GELU activation over the CIFAR-10 dataset for our future experiments.
Table 2: Clean and APGD accuracy of ResNet-18 model with different activation functions over CIFAR-10 test set before (Base) and after integration with MeanSparse technique for different value of $\alpha$
Activation
@@
A.2 Threat Model
In this experiment, we focused on the effectiveness of MeanSparse against two well-studied threat models, $\ell_{\infty}$ and $\ell_{2}$. Until now, our focus was on the $\ell_{\infty}$ threat model, which generally results in lower robust accuracy [9]. On the other hand, $\ell_{2}$ threat models can focus on a specific part of the image, and this localization of adversarial perturbation may challenge the effectiveness of the MeanSparse technique. In this experiment, we compare the robustness against both $\ell_{\infty}$ and $\ell_{2}$ threat models. The model is ResNet-18 with the GELU activation function as in our Activation Function experiment in 2. Table 3 illustrates the clean and APGD accuracy for $\ell_{\infty}$ and $\ell_{2}$ threats. As the accuracy for the $\ell_{\infty}$ threat model increases with increasing $\alpha$, the accuracy for $\ell_{2}$ also increases, which depicts the effectiveness and generalization of MeanSparse across well-studied threat models. Additionally, the improvement compared to the base model at $\alpha=0.2$ is larger for the $\ell_{2}$ threat than for the $\ell_{\infty}$ threat.
Table 3: Clean and APGD accuracy of ResNet-18 model with GELU activation function over CIFAR-10 test set before (Base) and after integration with MeanSparse technique for different value of $\alpha$ and different threat models from RobustBesnch [10]
Threat Model
@@ A.2 Threat Model
A.3 Blackbox Attacks
The MeanSparse operation masks the gradient during backward propagation when the input $a^{\mathrm{in}}$ lies within $\alpha\sigma_{\mathrm{ch}}$ of the channel mean $\mu_{\mathrm{ch}}$ (see the bottom plot in the first column of Figure 1). An interesting direction to explore is the performance of black-box attacks that do not rely on gradients, both before and after integrating MeanSparse. For this experiment, we select the Square Attack [2]. Table 4 shows the clean and robust accuracy for a ResNet-18 model with GELU activation before (Base) and after integration with MeanSparse (robust accuracy is denoted by Square, representing the accuracy under Square Attack). The results indicate a similar improvement in robust accuracy on par with APGD, while clean accuracy remains largely unaffected across a wide range of $\alpha$ values, from $0.05$ to $0.2$.
Table 4: Clean and Square accuracy (Square accuracy in the Square column represents the accuracy after applying square attack [2]) of ResNet-18 model with GELU activation function over CIFAR-10 test set before (Base) and after integration with MeanSparse technique for different value of $\alpha$
@@ A.3 Blackbox Attacks
A.4 Adaptive Attack
In this experiment, we investigate the potential for the attacker to enhance the attack by utilizing the knowledge that MeanSparse is employed to robustify the model. We propose a formulation to adapt the original PGD attack, incorporating the integration of MeanSparse to create a more powerful attack. In this scenario, we assume the attacker has complete access to the locations, mean, and variance vectors of the integrated MeanSparse operator throughout the model. Let $\bm{x}$ represent the input image and $\{\bm{z}_{i}\}$ the corresponding representations at the input of the MeanSparse modules in the model. When a perturbation $\bm{\delta}$ is added to the input ($\bm{x}+\bm{\delta}$), it results in new representations $\{\bm{z}_{i}^{\delta}\}$. In the adaptive PGD attack, we solve the following optimization problem to determine the perturbation $\bm{\delta}$:
$\operatorname*{arg\,max}_{\bm{\delta}}\quad\mathrm{CE}(f(\bm{x}+\bm{\delta}),y)-\lambda\sum_{i}\|\bm{m}_{i}\odot(\bm{z}_{i}-\bm{z}_{i}^{\delta})\|_{p}^{q}$
$\text{subject to}\quad\|\bm{\delta}\|_{\infty}\leq\epsilon,\quad\bm{x}+\bm{\delta}\in[0,1]^{D}$ (9)
where $\mathrm{CE}(\cdot,\cdot)$ is the cross-entropy loss, $f$ is the model, $y$ is the true label, $\lambda$ is a hyper-parameter, $i$ indexes different MeanSparse modules in the model, $\odot$ is the element-wise multiplication operator, $\epsilon$ controls the perturbation imperceptibility via the $\ell_{\infty}$ norm, and $D$ is the dimension of the input image. $\bm{m}_{i}$ represents a mask with the same dimensions as $\bm{z}_{i}$, identifying the locations in $\bm{z}_{i}$ affected by the MeanSparse module. For a given $\bm{z}$, the corresponding $\bm{m}$ can be computed element-wise as:

### v3

@@ IV Experiments
(a) CIFAR-10, $\ell_{\infty}=\frac{8}{255}$
(b) CIFAR-10, $\ell_{2}=0.5$
(c) ImageNet , $\ell_{\infty}=\frac{4}{255}$
(d) CIFAR-100, $\ell_{\infty}=\frac{8}{255}$
Figure 3: Original models performance along with their performance after integrating with MeanSparse technique. For CIFAR-10 dataset with $\ell_{\infty}$ attack, we have WideResNet-94-16 (Rank 1) [49], RaWideResNet-70-16 (Rank 4) [15] and WideResNet-70-16 (Rank 5) [50] while for $\ell_{2}$ attack we have WideResNet-70-16 (Rank 1) [50]. For ImageNet dataset with $\ell_{\infty}$ attack, we have Swin-L (Rank 2) [51], ConvNeXt-L (Rank 4) [52] and RaWideResNet (Rank 12) [15]. For CIFAR-100 dataset with $\ell_{\infty}$ attack, we have WideResNet-70-16 (Rank 1) [50] (all the rankings are based on RobustBench [16])
IV-A Evaluation Metrics
Throughout the experiments, we use Clean to represent the model accuracy for clean data. $\text{A-PGD}_{\text{ce}}$ is used as a robustness measure and represents accuracy after applying Auto PGD with Cross-Entropy (CE) loss. A-PGD is used as another robustness measure and represents accuracy after applying Auto-PGD with both CE and Difference of Logits Ratio (DLR) loss metrics. Finally, we use AA to represent AutoAttack accuracy. All the robustness metrics can be found in [53]. In our experiments, MeanSparse is applied post-training, eliminating the randomness typically introduced during training. Moreover, results from RobustBench for $\text{A-PGD}_{\text{ce}}$, A-PGD, and AA accuracies are highly consistent, showing negligible variability. Consequently, statistical significance tests are almost zero for all the results reported in this section.
@@ IV-A Evaluation Metrics
TABLE I: Clean accuracy and accuracy under A-PGD-CE attack [53] (represented by $\text{APGD}_{\text{ce}}$) for leading robust models on the CIFAR-10 and ImageNet-1k datasets (Base row) and after their integration with the MeanSparse technique (RaWideResNet [15] and WideResNet [50] are leading models in Robustbench ranking for CIFAR-10 dataset and ConvNeXt-L [52] and RaWideResNet [15] are highly ranked models for ImageNet dataset in Robustbench [16]). The best result in each column is in bold and the threshold selected through the proposed pipeline is in blue.
@@
IV-B Integration to State-of-the-Art Models
Figure 3 shows that integrating MeanSparse into the leading robust models in the RobustBench ranking [53] across various benchmark datasets achieves a new SOTA in AutoAttack accuracy. Figure 3(a) depicts the results for the Clean and AA metrics of the standard and MeanSparse-integrated versions of the WideResNet-94-16 [49], RaWideResNet-70-16 [15] and WideResNet-70-16 [50] over CIFAR-10 dataset. The original WideResNet-94-16 [49] currently holds the SOTA performance in the RobustBench ranking [53] for $\ell_{\infty}$ untargeted attack. After post-processing with the proposed pipeline, the Clean accuracy drops slightly (from $93.68\%$ to $93.63\%$), while the AA significantly increases (from $73.71\%$ to $75.28\%$), setting a new record for robustness in terms of AA on the CIFAR-10 dataset. Similar results can be obtained for RaWideResNet-70-16 [15] and WideResNet-70-16 [50]. Also, as a result of post-processing the robust model based on the WideResNet-70-16 architecture, the resulting AA accuracy ($71.41\%$) exceeds this value for the robust model based on the RaWideResNet architecture ($71.07\%$), even though WideResNet has a less complex architecture compared to RaWideResNet, which is equipped with the Squeeze-and-Excitation module.
Figure 3(b) represents the result of integrating MeanSparse into the current SOTA robust model (WideResNet-70-16 [50]) in terms of AA accuracy with $\ell_{2}$ distance metric. Similarly, while the clean accuracy is almost unchanged (from $95.54\%$ to $95.49\%$), the AA accuracy is improved (from $84.97\%$ to $87.28\%$).
Figure 3(c) illustrates the clean and AA accuracy for three different models adversarially trained on the ImageNet dataset (Swin-L [51], ConvNeXt-L [52] and RaWideResNet [15]). By carefully selecting the $\alpha$ value, while the Clean accuracy remains unchanged, the AA accuracy improves significantly. The Swin-L model [51] is currently the second-rank model on the ImageNet dataset. When integrating this model with MeanSparse, while clean accuracy drops slightly (from $78.92\%$ to $78.86\%$), the AA accuracy is increased (from $59.56\%$ to $62.12\%$) setting a new SOTA based on AA accuracy.
Figure 3(d) shows the results of integrating MeanSparse with the current state-of-the-art (SOTA) robust model on the CIFAR-100 dataset (WideResNet-70-16 [50]). For this model, the clean and AutoAttack (AA) accuracies are $75.22\%$ and $42.67\%$, respectively. After integrating MeanSparse, these accuracies change to $75.17\%$ and $44.78\%$, establishing a new SOTA robust model with nearly identical clean accuracy.
Table I presents the results of integrating MeanSparse into RaWideResNet [15] and WideResNet [50] architectures trained on the CIFAR-10 dataset and ConvNeXt-L [52] and RaWideResNet [15] on the ImageNet dataset using adversarial training. The results are provided for different values of $\alpha$ (note that the threshold equals $\alpha\times\sigma_{\mathrm{ch}}$). The experiment results over the CIFAR-10 test set depict that for a range of $\alpha$ below $0.25$, the clean accuracy has decreased negligibly or almost unchanged while $\text{APGD}_{\text{ce}}$is increasing. Although these results may be attributed to overfitting, the result over the CIFAR-10 training set also represents a similar trend. So an important takeaway from CIFAR-10 results in Table I is that the variations around the mean value of representation can be considered as a capacity provided by the model for the attackers while its utilization is limited. MeanSparse requires one pass over CIFAR-10 training set to calculate the statistics. The time for this pass is 102 seconds and 95 seconds for the RaWideResNet and WideResNet architectures, respectively. The increase in $\text{APGD}_{\text{ce}}$evaluation time after MeanSparse integration is negligible, with only a $2\%$ increase in time.
Results over ImageNet presented in Table I show that similar outcomes are achieved when MeanSparse is integrated into models based on the ConvNeXt-L [52] and RaWideResNet [15] architectures adversarially trained. Note that throughout our experiment over the RaWideResNet [15] architecture for both CIFAR-10 and ImageNet-1k datasets, we do not apply the sparsification operator inside the Squeeze-and-Excitation module. The time for the MeanSparse pass over the ImageNet-1K training set is 89 minutes for the ConvNeXt-L architecture and 101 minutes for the RaWideResNet architecture. The increase in $\text{APGD}_{\text{ce}}$evaluation time after MeanSparse integration is also negligible, with only a $3\%$ increase in time.
IV-C Activation Function
MeanSparse is generally appended before the activations function. This questions the effectiveness of MeanSparse for different activation functions. In this experiment, we check the performance of MeanSparse when integrating the models with different activation functions. The experiment trains a ResNet-18 model [32] on the CIFAR-10 dataset [47] for 200 epochs using SGD (initial learning rate: 0.1, weight decay: 0.0005, momentum: 0.9) with a batch size of 256. The learning rate is reduced by a factor of 10 at epochs 100 and 150. Adversarial training is performed using 10-step PGD [1] with respect to $\ell_{\infty}$ attacks, having a radius of $8/255$ and a step size of 0.0078. The best model is selected based on the highest PGD adversarial accuracy on the test set evaluated at each epoch.
@@ IV-C Activation Function
Table II demonstrates the Clean and APGD accuracy metrics for a base ResNet-18 model, denoted by Base, with different activation functions and the same model post-processed by the proposed pipeline to integrate MeanSaprse. We use both non-parametric and parametric activation functions, such as PSSiLU [9]. In our experiment, the GELU activation function, proposed in [36] and generally used in attention-based architectures, results in the best APGD accuracy. When this model is integrated with MeanSparse, we observe a significant improvement in APGD accuracy ($49.12\%\rightarrow 50.37\%$) while the clean accuracy slightly decreases ($84.59\%\rightarrow 84.32\%$) for $\alpha=0.2$. Additionally, if we can tolerate $\sim 1\%$ reduction in clean accuracy, we can increase APGD accuracy by more than $2\%$. The first important takeaway is the effectiveness of MeanSparse integration with different activations and the inability of the activation function to replicate the role of the MeanSparse technique. The second is that $\alpha=0.2$ consistently leads to large improvements in APDG accuracy with a negligible reduction in clean accuracy for all activation functions. Even for the ReLU activation function, $\alpha=0.2$ increases both Clean and APGD accuracies simultaneously. We will use the ResNet-18 model trained adversarially with GELU activation over the CIFAR-10 dataset for our future experiments.
TABLE II: Clean and APGD accuracy of ResNet-18 model with different activation functions over CIFAR-10 test set before (Base) and after integration with MeanSparse technique for different value of $\alpha$
@@
TABLE III: Clean and APGD accuracy of ResNet-18 model with GELU activation function over CIFAR-10 test set before (Base) and after integration with MeanSparse technique for different value of $\alpha$ and different threat models from RobustBesnch [53]
@@
IV-D Blackbox Attacks
The MeanSparse operation masks the gradient during backward propagation when the input $a^{\mathrm{in}}$ lies within $\alpha\sigma_{\mathrm{ch}}$ of the channel mean $\mu_{\mathrm{ch}}$ (see the bottom plot in the first column of Figure 1). An interesting direction to explore is the performance of black-box attacks that do not rely on gradients, both before and after integrating MeanSparse. For this experiment, we select the Square Attack [54]. Table IV shows the clean and robust accuracy for a ResNet-18 model with GELU activation before (Base) and after integration with MeanSparse (robust accuracy is denoted by Square, representing the accuracy under Square Attack). The results indicate a similar improvement in robust accuracy on par with APGD, while clean accuracy remains largely unaffected across a wide range of $\alpha$ values, from $0.05$ to $0.2$.
TABLE IV: Clean and Square accuracy (Square accuracy in the Square column represents the accuracy after applying square attack [54]) of ResNet-18 model with GELU activation function over CIFAR-10 test set before (Base) and after integration with MeanSparse technique for different value of $\alpha$
@@ IV-D Blackbox Attacks
IV-E Adaptive Attack
In this experiment, we investigate the potential for the attacker to enhance the attack by utilizing the knowledge that MeanSparse is employed to robustify the model. We propose a formulation to adapt the original PGD attack, incorporating the integration of MeanSparse to create a more powerful attack. In this scenario, we assume the attacker has complete access to the locations, mean, and variance vectors of the integrated MeanSparse operator throughout the model. Let $\boldsymbol{x}$ represent the input image and $\{\boldsymbol{z}_{i}\}$ the corresponding representations at the input of the MeanSparse modules in the model. When a perturbation $\boldsymbol{\delta}$ is added to the input ($\boldsymbol{x}+\boldsymbol{\delta}$), it results in new representations $\{\boldsymbol{z}_{i}^{\delta}\}$. In the adaptive PGD attack, we solve the following optimization problem to determine the perturbation $\boldsymbol{\delta}$:
$\operatornamewithlimits{arg\,max}_{\boldsymbol{\delta}}\quad\mathrm{CE}(f(\boldsymbol{x}+\boldsymbol{\delta}),y)-\lambda\sum_{i}\|\boldsymbol{m}_{i}\odot(\boldsymbol{z}_{i}-\boldsymbol{z}_{i}^{\delta})\|_{p}^{q}$
$\text{subject to}\quad\|\boldsymbol{\delta}\|_{\infty}\leq\epsilon,\quad\boldsymbol{x}+\boldsymbol{\delta}\in[0,1]^{D}$ (8)
where $\mathrm{CE}(\cdot,\cdot)$ is the cross-entropy loss, $f$ is the model, $y$ is the true label, $\lambda$ is a hyper-parameter, $i$ indexes different MeanSparse modules in the model, $\odot$ is the element-wise multiplication operator, $\epsilon$ controls the perturbation imperceptibility via the $\ell_{\infty}$ norm, and $D$ is the dimension of the input image. $\boldsymbol{m}_{i}$ represents a mask with the same dimensions as $\boldsymbol{z}_{i}$, identifying the locations in $\boldsymbol{z}_{i}$ affected by the MeanSparse module. For a given $\boldsymbol{z}$, the corresponding $\boldsymbol{m}$ can be computed element-wise as: